OpenStudy (anonymous):

solve: (x-3)^2=16 a) {-19,13} b) {-13,19} c) {-7,1} d) {-1,7}

OpenStudy (anonymous):

taking root of both sides, x-3 = +/- 4 so x = +4 +3 = 7 or x = -4 + 3 = -1 hence d) is the correct answer

OpenStudy (anonymous):

thanks so much dude. i was wayy off in my work

OpenStudy (anonymous):

if u r satisfied, pls click on "Good Answer" button so that the question is shown as solved

OpenStudy (anonymous):

thx :)